Recent Log Entries

COMMENT ON *Vacuum Activities CEBAF 03/03/20 through

Log Number: 3858951Logbook(s): Body: 

rather large-conductance leak found on CM03 during WG
removal. Pressure dropped because of atmospheric change ~
9Torr over an hour after the removal, unknown which window
caused the leak - luckily the pressure went down in the CM
rather than the other way round.

COMMENT ON Install PSS jumpers for laser work at the Injector

Log Number: 3858946Logbook(s): Body: 

In addition to previously listed signals, which will be bypassed, SSG will have to temporarily remove two PSS Laser Shutters from laser's path on the Laser Table. It was deemed impractical due to ongoing Injector segment upgrade work to attempt to power up these Shutters. Therefore these shutters will be moved away but their wiring will remain intact.

COMMENT ON Injector Upgrade - Phase 1 Installation (Gun to Chopper)

Log Number: 3858938Logbook(s): Body: 

Day #50, Week 11 Day 2, Plans:

Tunnel --
* MechInstall bringing in new Y-chamber and solenoid from
TL
* MechInstall moving gun table to new base pads
* CIS providing vacuum support for tunnel work

InjSB --
* Gun HV interlock work, IN01B{02-05} (Windham)
* Viewer cable terminations, IN03B07 (Norris)
* PSS Upgrades, IN02B{NI,29,30} +briefly elsewhere,
afternoons (Beizer)

Test Lab --
* V-wien HV testing on-going up to +/-20kV at UITF
(Hernandez-Garcia)
* New prebuncher with RF group for testing/tuning
(Wissman)
* A1/A2 cube assembly (Adderley)

MagTest --
* MHD coil (same on A1/A2 aperture

COMMENT ON Check and Adjust Current and Voltage readbacks, locally and remotely on all Legacy CPS.

Log Number: 3858910Logbook(s): Body: 

Calibration of the high voltage divider boxes located in zones 1L05 and 1L07 were completed. Calibration of the voltage and analog read backs as well as calibration of the current overload trip points in zones 1L13 and 1L07 were also completed. Finally, the current analog and digital read back for zone 1L07 was completed as well.

COMMENT ON Injector Upgrade - Phase 1 Installation (Gun to Chopper)

Log Number: 3858909Logbook(s): Body: 

Day #49, Week 11 Day 1, Wrap-Up:

AIPINJ --
* UITF Matasusada PSs powered up and HV enabled in
preparation for CEBAF H-wien HV evaluation. 3858862
* Y-chamber placed on carts in TL in preparation of move
to tunnel tomorrow

Other Injector SAD Work --
* PSS upgrades continue

VBVMA01 and VIPMA01 disabled in advance of target swap

Log Number: 3858907Logbook(s): Entry Makers: poelkerBody: 

The gate valve at the IBC (VBVMA01) has been disabled. It cannot be opened via epics command. ANd I unplugged the ion pump in the short region downstream of this valve and the IBC, because they plan to vent either side of the IBC tonight in preparation for target swap.

sharp current drops as reported at HDIce dump

Log Number: 3858905Logbook(s): Entry Makers: poelkerBody: 

We remain puzzled by the recuring dump current drop outs that happen with or without the dump bias applied. THese sharp current drops are sometimes correlated with temp increase at the He3/He4 mixing chamber. The target halo monitors don't obviously show a correlation with beam position or detector rates.

Figure 1 blue trace shows the dump current at about 840pA, which would correspond to ~ 1nA at the target. Beam current sharply dropping to 30% of desired value and then recovering over ~ 1 minute. There are 7 of these spikes in the figure.

COMMENT ON PSS North Linac Segment PLC Upgrade

Log Number: 3858903Logbook(s): Body: 

Worked in rack NA01B03, today removing old Red striped hookup wires, used in old PLC signal configuration. Begin the process of relocating existing field signal cables, that will be used with new PLC system upgrade. Work progressing.

1nA run HDIce

Log Number: 3858902Logbook(s): Entry Makers: poelkerBody: 

1 nA on the HDIce target. The striptool shows 840pA, but there's 1nA at the target, some beam does not make it to the dump. Tsuneo will tally data with DAQ using polarized e-beam (now) and unpolarized e-beam (turn OFF pockel cell)

COMMENT ON *Vacuum Activities CEBAF 03/03/20 through

Log Number: 3858898Logbook(s): Body: 

1L02 Installed upstream beampipe o-ring/insvac. Closed up access ports and reestablished vacuum. Rough leak check looks okay.

1L03 Preparing for ceramic warm window installations tomorrow.

1L04 & 1L06. Final insulating vacuum leak check completed. Insulating vacuum pumped for 48 hours. New upstream and downstream beampipe / insulating vacuum o-rings installed. Minimal detectable leak 1e-9. Access port o-rings looked really good. No sign of hardening or cracks.

iocsoftfac restored

Log Number: 3858890Logbook(s): Tags: Body: restorelib 2.2.1 Target restore date was 2020-11-30 14:54:10 The output from the myRestore execution was as follows: iocsoftfac DemandPower@2020-11-30 14:54 1(1) restored system@2020-11-30 14:54 2(2) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocsoftfac (success) Procedure return status: Success

two currents, two beam orbits? not likely

Log Number: 3858889Logbook(s): Entry Makers: poelkerBody: 

keV BPM wiresums at two currents, 1 and 0.5 uA for Tsuneo's DAQ test. Figure1 makes sense. Twice as much signal at 1uA compared to 0.5 uA.

But why are the beam positions different? Figure 2. I don't think the beam position actually changes.

Pages